### v1

# Schatten classes on noncommutative tori: Kernel conditions

Abstract.
In this note, we give criteria on noncommutative integral kernels ensuring that integral operators on quantum torus belong to Schatten classes. With the engagement of a noncommutative Schwartz’ kernel theorem on the quantum torus, a specific test for Schatten class properties of bounded operators on the quantum torus is established.
2000 Mathematics Subject Classification: Primary: 46L52, 46L51, 46L87. Secondary: 47L25, 43A99
Key words: Quantum tori, Schatten classes
1. Introduction
Given a closed smooth manifold, the properties of the integral operators defined on it merit thorough examination. For example, finding sufficient conditions of Schwartz integral kernels to ensure the corresponding integral operators belong to different Schatten classes. Identifying such criteria in various domains is a classical problem that has been extensively scrutinized. Specifically, it is well-established that the regularity and smoothness of the kernel are intrinsically linked to the asymptotic behaviors of the singular values.
In the paper by Delgado and Ruzhansky [2], one presented criteria for Schatten classes of integral operators on compact smooth manifolds, their approach is based on the renowned factorization technique, particularly in the manner elucidated by O’Brien [7]. Their results encompass compact Lie groups as a special case. The sufficient conditions for integral kernels $K(x,y)$ to belong to Schatten classes will necessitate regularity of a specific order in either $x$, $y$, or both. Their method’s advantage lies in the flexibility afforded to the sets of variables upon which the kernel’s regularity is imposed. Before long the authors significantly extended their main results of [2] to Euclidean space $\mathbb{R}^{n}$ by means of the anharmonic oscillators in [3]. The main results of [3] provide a non-compact counterpart of the results in [2]. Determining the anharmonic oscillators in the noncommutative setting is still an open question. Namely, to fill this gap if we consider the noncommutative Euclidean space and also establish the criteria for Schatten classes on it, but this will be done elsewhere.
Motivated by the work of Delgado and Ruzhansky [2], we aim to extend the aforementioned criteria for Schatten classes to the noncommutative setting. The inaugural fully non-commutative singular integral operator theory was given by González-Pérez, Junges, and Parcet [4]. Put differently, the very first form of an integral operator acting on a general von Neumann algebra was articulated by the three authors. A crucial point is to define kernels over ¯${\mathcal{M}}\overline{\otimes}({\mathcal{M}})^{\rm op}$ where the second copy is the opposite algebra of the von Neumann algebra ${\mathcal{M}}$.
As a particular instance of the aforesaid definition, we choose the quantum tori ${\mathbb{T}}_{\theta}^{d}$ to work with and establish the criteria for Schatten classes on it. Quantum tori (also known as noncommutative tori and irrational rotation algebras) are landmark examples in noncommutative geometry and can be viewed as an analog of closed manifolds in the non commutative setting. Many aspects of harmonic analysis have been studied for quantum tori and were taken as fundamental examples in noncommutative geometry and noncommutative analysis, see [1, 13, 6].
In this note, we add some regularity conditions on the noncommutative kernel $k\in L^{2}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$ by engaging the Bessel potentials on ${\mathbb{T}}_{\theta}^{d}$. The regularity of kernel $k$ will give us the Schatten class property of the associated integral operator $T_{k}$. And we consider a bounded linear operator $T$ on $L^{2}({\mathbb{T}}_{\theta}^{d})$, we first exploit a noncommutative Schwartz’ kernel theorem for $T$ in section 3.2, and then invoke the criteria we obtain in section 3.3 to give a characterization of Schatten classes for $T$. Our approach is adapted from [2] and [11].
2. Notation
2.1. Noncommutative Tori
Harmonic analysis on noncommutative tori is an established subject. The exposition here closely follows [13], and for the sake of brevity, we refer the reader to [13] for a detailed exposition of the topic and provide here only the definitions relevant to this text.
2.1.1. Basic definitions
Let $d\geq 2$ and let $\theta=(\theta_{kj})$ be a real skew symmetric $d\times d$-matrix. The associated $d$-dimensional noncommutative torus $\mathcal{A}_{\theta}$ is the universal $C^{*}$-algebra generated by $d$ unitary operators $U_{1},\ldots,U_{d}$ satisfying the following commutation relation
(2.1) $U_{k}U_{j}=e^{2\pi\mathrm{i}\theta_{kj}}U_{j}U_{k},\quad j,k=1,\ldots,d.$
We will use standard notation from multiple Fourier series. Let $U=(U_{1},\cdots,U_{d})$. For $m=(m_{1},\cdots,m_{d})\in{\mathbb{Z}}^{d}$ we define
$U^{m}=U_{1}^{m_{1}}\cdots U_{d}^{m_{d}}.$
A polynomial in $U$ is a finite sum
$x=\sum_{m\in\mathbb{Z}^{d}}\alpha_{m}U^{m}\quad\text{with}\quad\alpha_{m}\in\mathbb{C},$
that is, $\alpha_{m}=0$ for all but finitely many indices $m\in\mathbb{Z}^{d}.$ The involution algebra $\mathcal{P}_{\theta}$ of all such polynomials is dense in ${\mathcal{A}}_{\theta}.$ For any polynomial $x$ as above we define
$\tau(x)=\alpha_{0},$
where 0=(0,,0)$0$=(0$\cdots$,0)0 = ( 0 , ⋯ , 0 ). Then, $\tau$ extends to a faithful tracial state on ${\mathcal{A}}_{\theta}$. Let $\mathbb{T}^{d}_{\theta}$ be the $w^{*}$-closure of ${\mathcal{A}}_{\theta}$ in the GNS representation of $\tau$. This is our $d$-dimensional quantum torus. The state $\tau$ extends to a normal faithful tracial state on $\mathbb{T}^{d}_{\theta}$ that will be denoted again by $\tau$. Recall that the von Neumann algebra $\mathbb{T}^{d}_{\theta}$ is hyperfinite.
The $L^{p}$-spaces for $p\in[1,\infty)$ on $\mathbb{T}^{d}_{\theta}$ are then defined as the operator $L^{p}$-spaces on $(L^{\infty}(\mathbb{T}^{d}_{\theta}),\tau)$,
$L^{p}(\mathbb{T}^{d}_{\theta}):=L^{p}(L^{\infty}(\mathbb{T}^{d}_{\theta}),\tau).$
Any $x\in L^{1}({\mathbb{T}}^{d}_{\theta})$ admits a formal Fourier series:
$x\sim\sum_{m\in\mathbb{Z}^{d}}\widehat{x}(m)U^{m},$
where
$\widehat{x}(m)=\tau((U^{m})^{*}x),\quad m\in\mathbb{Z}^{d}$
are the Fourier coefficients of $x$. The operator $x$ is, of course, uniquely determined by its Fourier series. For $x\in L^{2}(\mathbb{T}^{d}_{\theta})$, standard Hilbert space arguments show that it can be written as an $L_{2}$-convergent series:
$x=\sum_{n\in{\mathbb{Z}}^{d}}\widehat{x}(n)U^{n},$
with
(2.2) $\|x\|_{2}^{2}=\sum_{n\in{\mathbb{Z}}^{d}}|\widehat{x}(n)|^{2}.$
The space $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ is defined to be the subset of $C({\mathbb{T}}_{\theta}^{d})$ with the rapid decay sequence of Fourier coefficients $\{\widehat{x}_{n}\}_{n\in\mathbb{Z}^{d}}$. There is a canonical Fréchet topology on $C^{\infty}({\mathbb{T}}_{\theta}^{d})$, the topological dual space $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$ of $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ is called the distribution space on ${\mathbb{T}}_{\theta}^{d}$.
2.1.2. Harmonic analysis on quantum tori
Many aspects of harmonic analysis on ${\mathbb{T}}^{d}$ carry foward to $\mathbb{T}^{d}_{\theta}$. In the following, we give the definitions of partial differentiations and Fourier multipliers.
Definition 2.1.
The partial differentiation operators $\partial_{j}$, $j=1,\cdots,d$ on $\mathbb{T}^{d}_{\theta}$ are defined as:
$\partial_{j}(U^{n})=2\pi{\rm{i}}n_{j}U^{n},\quad n=(n_{1},\ldots,n_{d})\in{\mathbb{Z}}^{d}.$
Every partial derivation $\partial_{j}$ can be viewed as a densely defined closed (unbounded) operator on $L^{2}(\mathbb{T}^{d}_{\theta})$, whose adjoint is equal to $-\partial_{j}$. Let $\Delta=\partial_{1}^{2}+\cdots+\partial_{d}^{2}$ be the Laplacian. Then $\Delta=-(\partial_{1}^{*}\partial_{1}+\cdots+\partial_{d}^{*}\partial_{d})$, so $-\Delta$ is a positive operator on $L^{2}(\mathbb{T}^{d}_{\theta})$ with spectrum equal to $\{4\pi^{2}|n|^{2}:n\in{\mathbb{Z}}^{d}\}$. As in the Euclidean case, we let $D_{j}=-{\rm{i}}\partial_{j}$, which is then self-adjoint. Given $n=(n_{1},\cdots,n_{d})\in\mathbb{N}_{0}^{d}$ ($\mathbb{N}_{0}$ denoting the set of nonnegative integers), the associated partial derivation $D^{n}$ is defined to be $D_{1}^{n_{1}}\cdots D_{d}^{n_{d}}$. The order of $D^{n}$ is $|n|_{1}=n_{1}+\cdots+n_{d}$. By duality, the derivations transfer to $\mathcal{D}^{\prime}(\mathbb{T}^{d}_{\theta})$ as well.
Definition 2.2.
Let $g$ be a bounded scalar function on ${\mathbb{Z}}^{d}$. For $x\in L_{2}({\mathbb{T}}_{\theta}^{d})$, the Fourier multiplier $T_{g}$ with symbol $g$ is defined on $x$ by:
(2.3) $T_{g}x=\sum_{n\in{\mathbb{Z}}^{d}}g(n)\widehat{x}(n)U^{n}.$
By virtue of the Plancherel identity (2.2), $T_{g}$ indeed defines a bounded linear operator on $L^{2}({\mathbb{T}}_{\theta}^{d})$ and the above series converges in the $L^{2}$-sense. If $g$ is unbounded, we may define $T_{g}$ on the dense subspace of $L^{2}({\mathbb{T}}_{\theta}^{d})$ of those $x$ with finitely many non-zero Fourier coefficients.
More generally, if $g$ is a bounded scalar function on ${\mathbb{R}}^{d}$, we use the same notation $T_{g}$ to denote the Fourier multiplier with the symbol $g|_{{\mathbb{Z}}^{d}}$. If $g$ is a bounded scalar function on ${\mathbb{R}}^{d}\setminus\{0\}$ (or ${\mathbb{Z}}^{d}\setminus\{0\}$), $T_{g}$ is then a Fourier multiplier on the subspace of $\mathcal{D}^{\prime}({\mathbb{T}}^{d}_{\theta})$ of all $x$ such that $\widehat{x}(0)=0$.
Typical examples of Fourier multipliers are quantum analogs of Bessel and Riesz potentials. Let $\alpha\in{\mathbb{R}}$. Define $J_{\alpha}$ on ${\mathbb{R}}^{d}$ and $I_{\alpha}$ on ${\mathbb{R}}^{d}\setminus\{0\}$ by
$J_{\alpha}(\xi)=(1+|\xi|^{2})^{\frac{\alpha}{2}}\;\text{ and }\;I_{\alpha}(\xi)=|\xi|^{\alpha}\,.$
Their associated Fourier multipliers are the Bessel and Riesz potentials of order $\alpha$, denoted by $J^{a}$ and $I^{\alpha}$, respectively. By duality, $J^{\alpha}$ is also a Fourier multiplier on $\mathcal{D}^{\prime}({\mathbb{T}}^{d}_{\theta})$, and $I^{\alpha}$ a Fourier multiplier on the subspace of $\mathcal{D}^{\prime}({\mathbb{T}}^{d}_{\theta})$ of all $x$ such that $\widehat{x}(0)=0$. Note that
$J^{\alpha}=(1-(2\pi)^{-2}\Delta)^{\frac{\alpha}{2}}\;\text{ and }\;I^{\alpha}=(-(2\pi)^{-2}\Delta)^{\frac{\alpha}{2}}.$
It is worthwhile adverting that since the function $n\mapsto(1+|n|^{2})^{-\frac{\alpha}{2}}$ belongs to $\ell_{\frac{d}{\alpha},\infty}$, for any $\alpha>0$ we have
$J^{-\alpha}\in S_{\frac{d}{\alpha},\infty}.$
Analogously, when $\alpha>\frac{d}{p}$, it suggests that
$J^{-\alpha}\in S_{p}.$
For $\alpha\in\mathbb{R}$, the potential Sobolev space of order $\alpha$ is defined to be
$H_{p}^{\alpha}({\mathbb{T}}_{\theta}^{d})=\{x\in\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d}):J^{\alpha}x\in L^{p}({\mathbb{T}}_{\theta}^{d})\},$
with the norm
$\|x\|_{H_{p}^{\alpha}({\mathbb{T}}_{\theta}^{d})}=\|J^{\alpha}x\|_{p}.$
We now define the Sobolev space with mixed regularity $\alpha_{1},\alpha_{2}\geq 0$.
Definition 2.3.
Let $k\in L^{2}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})$ and $\alpha_{1},\alpha_{2}\geq 0$, we say that $k\in H^{\alpha_{1},\alpha_{2}}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$ if and only if $(J^{\alpha_{1}}\otimes J^{\alpha_{2}})k\in L^{2}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$.
2.2. Schatten classes and Schur multipliers
In this subsection let us briefly collect material concerning Schatten classes; for more details, we refer the reader to [12, 8, 9, 5]. Let $H$ be a complex separable Hilbert space, and let ${\mathcal{B}}(H)$ denote the set of bounded operators on $H$, and let ${\mathcal{K}}(H)$ denote the ideal of compact operators on $H$. Given $T\in{\mathcal{K}}(H)$, the sequence of singular values $\mu(T)=\{\mu(k,T)\}_{k=0}^{\infty}$ is defined as
$\mu(k,T)=\inf\{\|T-R\|\;:\;\mathrm{rank}(R)\leq k\}.$
Equivalently, $\mu(T)$ is the sequence of eigenvalues of $|T|$ arranged in non-increasing order with multiplicities.
Let $p\in(0,\infty).$ The Schatten class $S_{p}$ is the set of operators $T$ in ${\mathcal{K}}(H)$ such that $\mu(T)$ is $p$-summable, i.e. in the sequence space $\ell_{p}$. If $p\geq 1$ then the $S_{p}$ norm is defined as
$\|T\|_{p}:=\|\mu(T)\|_{\ell_{p}}=\Big{(}\sum_{k=0}^{\infty}\mu(k,T)^{p}\Big{)}^{1/p}.$
With this norm, $S_{p}$ is a Banach space, and an ideal of ${\mathcal{B}}(H)$. For $p=\infty$ we set $S_{\infty}={\mathcal{B}}(H)$ equipped with the operator norm.
The weak Schatten class $S_{p,\infty}$ is defined in an analogous way, by setting operators $T$ such that $\mu(T)$ is in the weak $L^{p}$ space $\ell_{p,\infty}$ with the quasi norm
$\|T\|_{p,\infty}=\sup_{k\geq 0}(k+1)^{1/p}\mu(k,T)<\infty.$
As with the $S_{p}$ spaces, $S_{p,\infty}$ is an ideal of ${\mathcal{B}}(H)$.
Another equivalent way of defining the above Schatten classes is to consider the whole algebra ${\mathcal{B}}(H)$ with the usual trace Tr; then the Schatten class $S_{p}$ is defined as the noncommutative $L_{p}$ spaces $L_{p}({\mathcal{B}}(H),\mbox{\rm Tr})$, see e.g. [10, 14].
3. Schatten properties of integral operators on noncommutative tori
3.1. Schwartz’ kernel theorem
In the subsequent discussion, we will show that for a continuous linear operator
$T:C^{\infty}({\mathbb{T}}_{\theta}^{d})\to\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$
there exists a kernel $k\in\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})\otimes\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$ such that
$\langle Tx,y\rangle=\langle T_{k}x,y\rangle=\langle k,y\otimes x\rangle.$
Consider the space $\mathcal{A}$ of all the separately bilinear functionals $A$ on $C^{\infty}({\mathbb{T}}_{\theta}^{d})\times C^{\infty}({\mathbb{T}}_{\theta}^{d})$. Any distribution $\omega\in\mathcal{D}^{\prime}({\mathbb{T}}^{d}_{\theta}\otimes{\mathbb{T}}^{d}_{\theta})\simeq\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ gives rise to such a functional by specialization to the tensor product of two elements:
(3.1) $\Lambda\omega(x,y):=\omega(x\otimes y):=A(x,y).$
The kernel theorem states that the mapping:
$\Lambda:\omega\mapsto A$
is a linear homomorphism between $\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ and $\mathcal{A}$.
Theorem 3.1.
For any continuous linear functional $A$ on $C^{\infty}({\mathbb{T}}_{\theta}^{d})\times C^{\infty}({\mathbb{T}}_{\theta}^{d})$ there exists precisely one distribution $\omega\in\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ such that
(3.2) $A(x,y)=\langle\omega,x\otimes y\rangle$
holds for all $(x,y)\in C^{\infty}({\mathbb{T}}_{\theta}^{d})\times C^{\infty}({\mathbb{T}}_{\theta}^{d})$. The mapping $A\mapsto\omega$ in (3.2) is a linear homeomorphism.
Proof.
Every element $h\in C^{\infty}({\mathbb{T}}^{d}_{\theta}\otimes{\mathbb{T}}^{d}_{\theta})$ can be written in the form
$h=\sum_{m,n\in\mathbb{Z}^{d}}b_{m,n}U^{m}\otimes U^{n},$
where
(3.3) $\begin{split}b_{m,n}&=\tau\otimes\tau(h(U^{m}\otimes U^{n})^{*})\\ &=\langle h,U^{m}\otimes U^{n}\rangle\\ &=\langle(J^{-\alpha_{1}}\otimes J^{-\alpha_{2}})(J^{\alpha_{1}}\otimes J^{\alpha_{2}})h,U^{m}\otimes U^{n}\rangle\\ &=\langle(J^{\alpha_{1}}\otimes J^{\alpha_{2}})h,(J^{-\alpha_{1}}\otimes J^{-\alpha_{2}})(U^{m}\otimes U^{n})\rangle.\end{split}$
One note that in the previous identities, the indices $\alpha_{1},\alpha_{2}$ are arbitrary, we replace them with $\alpha_{1}+s_{0},\alpha_{2}+s_{0}$ respectively, where $s_{0}\in\mathbb{R}$ is a constant. The Cauchy-Schwartz inequality yields that
(3.4) $\begin{split}|b_{m,n}|&=|\langle(J^{\alpha_{1}+s_{0}}\otimes J^{\alpha_{2}+s_{0}})h,(J^{-(\alpha_{1}+s_{0})}\otimes J^{-(\alpha_{2}+s_{0})})(U^{m}\otimes U^{n})\rangle|\\ &\leq\|(J^{\alpha_{1}+s_{0}}\otimes J^{\alpha_{2}+s_{0}})h\|_{L^{2}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})}\|(J^{-(\alpha_{1}+s_{0})}\otimes J^{-(\alpha_{2}+s_{0})})(U^{m}\otimes U^{n})\|_{L^{2}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})}\\ &=\|(J^{\alpha_{1}+s_{0}}\otimes J^{\alpha_{2}+s_{0}})h\|_{L^{2}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})}\|(1+|m|^{2})^{-\frac{\alpha_{1}+s_{0}}{2}}(1+|n|^{2})^{-\frac{\alpha_{2}+s_{0}}{2}}(U^{m}\otimes U^{n})\|_{L^{2}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})}\\ &\lesssim\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}}(1+|m|^{2})^{-\frac{\alpha_{1}+s_{0}}{2}}(1+|n|^{2})^{-\frac{\alpha_{2}+s_{0}}{2}},\end{split}$
and choose $s_{0}$ to be an integer such that $\sum_{m\in\mathbb{Z}^{d}}(1+|m|^{2})^{-\frac{s_{0}}{2}}<\infty$.
We can rewrite $h$ as
(3.5) $\begin{split}h=&\sum_{m,n\in\mathbb{Z}^{d}}(1+|m|^{2})^{\frac{\alpha_{1}}{2}}(1+|n|^{2})^{\frac{\alpha_{2}}{2}}b_{m,n}((1+|m|^{2})^{-\frac{\alpha_{1}}{2}}U^{m}\otimes(1+|n|^{2})^{-\frac{\alpha_{2}}{2}}U^{n})\\ &=\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}V^{m}\otimes V^{n},\end{split}$
where $\alpha_{1},\alpha_{2}\geq 0$ and
(3.6) $\begin{split}a_{m,n}&=(1+|m|^{2})^{\frac{\alpha_{1}}{2}}(1+|n|^{2})^{\frac{\alpha_{2}}{2}}b_{m,n},\\ V^{m}&=(1+|m|^{2})^{-\frac{\alpha_{1}}{2}}U^{m},\ V^{n}=(1+|n|^{2})^{-\frac{\alpha_{2}}{2}}U^{n}.\end{split}$
Note that
$|a_{m,n}|\lesssim\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}}(1+|m|^{2})^{-\frac{s_{0}}{2}}(1+|n|^{2})^{-\frac{s_{0}}{2}}$
and
$\|V^{m}\|\leq 1,\ \|V^{n}\|\leq 1$
since $\alpha_{1},\alpha_{2}\geq 0$.
We are in a position to construct $\omega$, define
$\langle\omega,h\rangle=\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n}),$
we conclude from (3.4) and (3.6) that
$\begin{split}|\langle\omega,h\rangle|&=|\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n})|\\ &\leq\sum_{m,n\in\mathbb{Z}^{d}}|a_{m,n}||A(V^{m},V^{n})|\\ &\lesssim\sum_{m,n\in\mathbb{Z}^{d}}\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}}(1+|m|^{2})^{-\frac{s_{0}}{2}}(1+|n|^{2})^{-\frac{s_{0}}{2}}\|A\|\\ &\leq C\|h\|_{H^{\alpha_{1}+s_{0},\alpha_{2}+s_{0}}},\end{split}$
where the constant $C$ depends on $A$ and $s_{0}$. Hence $\omega$ is a distribution on ${\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}$. It is apparently the case that $\omega$ satisfies (3.2) and $\omega$ is uniquely determined by $A$ since the set of all the finite sums $\sum_{m,n\in\mathbb{Z}^{d}}b_{m,n}U^{m}\otimes U^{n}$ is dense in $C^{\infty}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})$.
Now we show that the map $\Lambda:\omega\mapsto A$ is a linear homeomorphism. To this end, we need to show that $\Lambda$ and $\Lambda^{-1}$ are both continuous.
The Fréchet topologies on $\mathcal{A}$ and $\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ can be expressed respectively as the following seminorms:
(3.7) $\begin{split}\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(A)&={\rm{sup}}\{|A(x,y)|,\ x,y\in B_{{\mathbb{T}}_{\theta}^{d}}\},\\ \rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega)&={\rm{sup}}\{|\langle\omega,h\rangle|,\,h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}\},\end{split}$
where $B_{{\mathbb{T}}_{\theta}^{d}}$ is a bounded set in $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ and $B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}$ is a bounded set in $C^{\infty}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})$.
Let $\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}$ be an arbitrary seminorm in $\mathcal{A}$. Then
(3.8) $\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(\Lambda\omega)=\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(A)=\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|A(x,y)|=\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|\langle\omega,x\otimes y\rangle|.$
It is clear that there exists a bounded set $B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}\subset C^{\infty}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})$ such that $B_{{\mathbb{T}}_{\theta}^{d}}\otimes B_{{\mathbb{T}}_{\theta}^{d}}\subset B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}$. Hence
$\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|\langle\omega,x\otimes y\rangle|\leq\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|\langle\omega,h\rangle|=\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega),$
which implies that $\Lambda$ is continuous.
Conversely, as for the expansion of $h$ in (3.5), since $h$ is bounded in $C^{\infty}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})$, we can restrict $V^{m},V^{n}$ to be bounded in $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ and $\sum_{m,n\in\mathbb{Z}^{d}}|a_{m,n}|\leq 1$. If $\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}$ is a seminorm on $C^{\infty}({\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d})$, we find
$\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\Lambda^{-1}A)=\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\omega)=\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|\langle\omega,h\rangle|=\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}A(V^{m},V^{n})|.$
Since $V^{m},V^{n}$ are bounded, we can find a bounded set $B_{{\mathbb{T}}_{\theta}^{d}}$ which contains $V^{m},V^{n}$. Thus,
$\rho_{B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}(\Lambda^{-1}A)\leq\sum_{m,n\in\mathbb{Z}^{d}}a_{m,n}\sup_{h\in B_{{\mathbb{T}}_{\theta}^{d}\otimes{\mathbb{T}}_{\theta}^{d}}}|A(V^{m},V^{n})|\leq\sup_{x,y\in{B_{{\mathbb{T}}_{\theta}^{d}}}}|A(x,y)|=\rho_{B_{{\mathbb{T}}_{\theta}^{d}},B_{{\mathbb{T}}_{\theta}^{d}}}(A),$
which demonstrates that $\Lambda^{-1}$ is continuous. ∎
Remark 3.2.
Leveraging the seminal transference method of Chen, Xu, and Yin in [1], we can transfer the problem of ${\mathbb{T}}_{\theta}^{d}$ to the corresponding ones of the operator-valued function in $L^{\infty}({\mathbb{T}}^{d},{\mathbb{T}}_{\theta}^{d})$. This transference method will provide an alternative proof of Theorem 3.1 by utilizing the operator-valued Schwartz’ kernel theorem, a subject that has yielded a wealth of results. However, the demonstration of the operator-valued theorem is as burdensome as directly proving it on the quantum tori. Moreover, the approach by transference is deficient in the discussions on the topology of $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$, rendering it somewhat superficial.
By consolidating the conclusions we have proven, we can ascertain that for any continuous linear operator $T:C^{\infty}({\mathbb{T}}_{\theta}^{d})\to\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$, there exists a Schwartz kernel $k\in\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ such that
$\langle Tx,y\rangle=\langle T_{k}x,y\rangle\ \text{for}\ x,y\in C^{\infty}({\mathbb{T}}_{\theta}^{d}),$
where $T_{k}$ is the operator associated to the kernel $k$ implemented by $\langle T_{k}x,y\rangle=\langle k,y\otimes x\rangle$ for $x,y\in C^{\infty}({\mathbb{T}}_{\theta}^{d})$.
Now for any $\theta$ and $1\leq p\leq\infty$, the space $L^{p}({\mathbb{T}}_{\theta}^{d})$ embeds into $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$: an element $x\in L^{p}({\mathbb{T}}_{\theta}^{d})$ derives a continuous functional on $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ by $y\mapsto\tau(xy)$. A map $T\in B(L^{2}({\mathbb{T}}_{\theta}^{d}))$ can be viewed as a linear map from $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ to $\mathcal{D}^{\prime}({\mathbb{T}}_{\theta}^{d})$, by plugging into Theorem 3.1, we obtain that there exists $k\in\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$ such that
$\langle Tx,y\rangle=\langle k,y\otimes x\rangle.$
Since $C^{\infty}({\mathbb{T}}_{\theta}^{d})$ is dense in $L^{2}({\mathbb{T}}_{\theta}^{d})$, the above identity holds for $x,y\in L^{2}({\mathbb{T}}_{\theta}^{d})$.
It is worth noting that the trace $\tau$ on ${\mathbb{T}}_{\theta}^{d}$ and $({\mathbb{T}}_{\theta}^{d})^{\rm{op}}$ coincides and the Bessel potential $J^{\alpha}$ acts in the same way for both spaces. So, we have that $L^{2}({\mathbb{T}}_{\theta}^{d})\simeq L^{2}(({\mathbb{T}}_{\theta}^{d})^{\rm op})$. Indeed, for the opposite algebra $({\mathbb{T}}_{\theta}^{d})^{\rm op}$, the opposite algebra structure gives the reversed product law for the elements in ${\mathbb{T}}_{\theta}^{d}$, namely, if we denote by $\star$ the product operation in $({\mathbb{T}}_{\theta}^{d})^{\rm op}$, then for the generating elements $U_{1},...,U_{d}$, we gain that
$U_{k}\star U_{j}=U_{j}U_{k}=e^{-2\pi\mathrm{i}\theta_{kj}}U_{k}U_{j}=e^{-2\pi\mathrm{i}\theta_{kj}}U_{j}\star U_{k},\quad j,k=1,\ldots,d.$
By the universal property of quantum tori, we see that $({\mathbb{T}}_{\theta}^{d})^{\rm{op}}$ is actually isomorphic to ${\mathbb{T}}_{-\theta}^{d}$.
If we assume that $k\in L^{2}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$, by the above clarifications, $L^{2}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})\subset\mathcal{D}^{\prime}({\mathbb{T}}^{2d}_{\theta\oplus\theta})$. And
(3.9) $\begin{split}\langle Tx,y\rangle&=\langle T_{k}x,y\rangle\\ &=\langle k,y\otimes x\rangle\\ &=\tau\otimes\tau(k(y\otimes x))\\ &=\tau(({\rm id}\otimes\tau)(k(1\otimes x))y),\end{split}$
which unveals that $Tx=T_{k}x=({\rm id}\otimes\tau)(k(1\otimes x))$ for $x\in L^{2}({\mathbb{T}}_{\theta}^{d})$.
3.2. Schatten classes on noncommutative tori
Now, appealing to (3.9), if we demand futhermore that $k\in H^{\alpha_{1},\alpha_{2}}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$, we can deduce that
$T\in S_{r}(L_{2}({\mathbb{T}}_{\theta}^{d}))\ \text{for}\ r>\frac{2d}{d+2(\alpha_{1}+\alpha_{2})}.$
The conclusion gives us a characterization of the Schatten class property for $T\in B(L^{2}({\mathbb{T}}_{\theta}^{d}))$ and is presented in the following theorem.
Theorem 3.3.
Let ${\mathbb{T}}_{\theta}^{d}$ be the quantum tori of dimension $d$ with $d\geq 2$ and let $a_{1},\alpha_{2}\geq 0$. Suppose $k\in H^{\alpha_{1},\alpha_{2}}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$. Then the integral operator $T_{k}$ on $L^{2}({\mathbb{T}}_{\theta}^{d})$ defined by
$T_{k}(x)=({\rm id}\otimes\tau)(k(1\otimes x))=({\rm id}\otimes\tau)((1\otimes x)k),$
is in the Schatten class $S_{r}(L^{2}({\mathbb{T}}_{\theta}^{d}))$ for
$r>\frac{2d}{d+2(\alpha_{1}+\alpha_{2})}.$
Proof.
Let $J^{\alpha}$ denote the Bessel potential of order $\alpha$ on quantum tori, then $J^{\alpha}$ can be viewed as a Fourier multiplier on $L^{2}({\mathbb{T}}_{\theta}^{d})$ with symbol $(1+|n|^{2})^{\alpha/2}$ for $n\in\mathbb{Z}^{d}$. And we get that $J^{{-\alpha_{2}}}\in S_{p_{2}}(L^{2}({\mathbb{T}}_{\theta}^{d}))$ if
$\alpha_{2}>\frac{d}{p_{2}}.$
If $x\in L^{2}({\mathbb{T}}_{\theta}^{d})$, we can write $x$ formally as $x=\sum_{n\in\mathbb{Z}^{d}}\widehat{x}(n)U^{n}$, which yields that
$\begin{split}J^{{-\alpha_{2}}}(x)&=\sum_{n\in\mathbb{Z}^{d}}(1+|n|^{2})^{\frac{-\alpha_{2}}{2}}\widehat{x}(n)U^{n}\\ &=\sum_{n\in\mathbb{Z}^{d}}(1+|n|^{2})^{\frac{-\alpha_{2}}{2}}\tau((U^{n})^{*}x)U^{n}\\ &=\sum_{n\in\mathbb{Z}^{d}}(1+|n|^{2})^{\frac{-\alpha_{2}}{2}}({\rm id}\otimes\tau)((U^{n}\otimes(U^{n})^{*})(1\otimes x))\\ &=({\rm id}\otimes\tau)(\sum_{n\in\mathbb{Z}^{d}}(1+|n|^{2})^{\frac{-\alpha_{2}}{2}}(U^{n}\otimes(U^{n})^{*})(1\otimes x)),\end{split}$
which means that $J^{-\alpha_{2}}(x)=T_{k_{2}}(x)$ for
$k_{2}=\sum_{n\in\mathbb{Z}^{d}}(1+|n|^{2})^{\frac{-\alpha_{2}}{2}}(U^{n}\otimes(U^{n})^{*}).$
Thus we have determined the kernel $k_{2}$ for the operator $J^{-\alpha_{2}}$ and $J^{\alpha_{2}}T_{k_{2}}(x)=x$.
We observe that for $x\in L^{2}({\mathbb{T}}_{\theta}^{d})$,
$\begin{split}\tau((U^{n})^{*}J^{\alpha_{1}}T_{k}(x))&=\tau((U^{n})^{*}J^{\alpha_{1}}({\rm id}\otimes\tau)(k(1\otimes x)))\\ &=\tau((U^{n})^{*}(J^{\alpha_{1}}\otimes\tau)(k(1\otimes x)))\\ &=\tau((U^{n})^{*}({\rm id}\otimes\tau)(J^{\alpha_{1}}\otimes{\rm id})(k(1\otimes x)))\\ &=\tau(({\rm id}\otimes\tau)((U^{n})^{*}\otimes 1)(J^{\alpha_{1}}\otimes{\rm id})(k(1\otimes x)))\\ &=\tau\otimes\tau(((U^{n})^{*}\otimes 1)(J^{\alpha_{1}}\otimes{\rm id})(k(1\otimes x))\\ &=\tau\otimes\tau(((U^{n})^{*}\otimes 1)(J^{\alpha_{1}}\otimes{\rm id})(k(1\otimes J^{\alpha_{2}}T_{k_{2}}(x))))\\ &=\tau\otimes\tau(((U^{n})^{*}\otimes 1)(J^{\alpha_{1}}\otimes{\rm id})k({\rm id}\otimes J^{\alpha_{2}})(1\otimes T_{k_{2}}(x)))\\ &=\langle((U^{n})^{*}\otimes 1)(J^{\alpha_{1}}\otimes{\rm id})k,({\rm id}\otimes J^{\alpha_{2}})(1\otimes T_{k_{2}}(x))^{*}\rangle\\ &=\langle({\rm id}\otimes J^{\alpha_{2}})((U^{n})^{*}\otimes 1)(J^{\alpha_{1}}\otimes{\rm id})k,(1\otimes T_{k_{2}}(x))^{*}\rangle\\ &=\langle((U^{n})^{*}\otimes 1)({\rm id}\otimes J^{\alpha_{2}})(J^{\alpha_{1}}\otimes{\rm id})k,(1\otimes T_{k_{2}}(x))^{*}\rangle\\ &=\tau\otimes\tau((U^{n})^{*}\otimes 1)(J^{\alpha_{1}}\otimes J^{\alpha_{2}})k(1\otimes T_{k_{2}}(x)))\\ &=\tau((U^{n})^{*}({\rm id}\otimes\tau)((J^{\alpha_{1}}\otimes J^{\alpha_{2}})k(1\otimes T_{k_{2}}(x))))\\ &=\tau((U^{n})^{*}T_{(J^{\alpha_{1}}\otimes J^{\alpha_{2}})k}(T_{k_{2}}x)),\end{split}$
[…]
Now we set $k_{1}=(J^{\alpha_{1}}\otimes J^{\alpha_{2}})k$, we have demonstrated that
$J^{\alpha_{1}}T_{k}=T_{k_{1}}T_{k_{2}}.$
By the assumption that $k\in H^{\alpha_{1},\alpha_{2}}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$ we know that $k_{1}\in L^{2}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$ and thus $T_{k_{1}}\in S_{2}$. With the restriction on $\alpha_{2}$, $T_{k_{2}}\in S_{p_{2}}$. By the Hölder inequality, we see that,
$J^{\alpha_{1}}T_{k}\in S_{t},$
with $\frac{1}{t}=\frac{1}{2}+\frac{1}{p_{2}}$.
On the other hand, since $J^{-\alpha_{1}}\in S_{p_{1}}$ when $\alpha_{1}>\frac{d}{p_{1}}$, we deduce that
$T_{k}=J^{-\alpha_{1}}J^{\alpha_{1}}T_{k}\in S_{r}$
with
$\frac{1}{r}=\frac{1}{t}+\frac{1}{p_{1}}=\frac{1}{2}+\frac{1}{p_{1}}+\frac{1}{p_{2}}.$
Engaging the inequalities
$\alpha_{1}>\frac{d}{p_{1}}\ \text{and}\ \alpha_{2}>\frac{d}{p_{2}},$
this is equivalent to saying that
$r>\frac{2d}{d+2(\alpha_{1}+\alpha_{2})}.$
Now, we conclude the proof by discussing the critical cases:
(1). If $\alpha_{1}=0$ and $\alpha_{2}>0$, just by eliminating the operator $J^{\alpha_{1}}$ in the above reasoning, the rest proceed in the same scheme,
$T_{k}\in S_{t},$
with $\frac{1}{t}=\frac{1}{2}+\frac{1}{p_{2}}$ and $\alpha_{2}>\frac{d}{p_{2}}$, we deduce that
$t>\frac{2d}{d+2\alpha_{2}}.$
(2). If $\alpha_{2}=0$ and $\alpha_{1}>0$, we proceed by taking the adjoint of the operator $T_{k}$, and by [4, Lemma 2.2], $T_{k}^{*}=T_{{\rm filp}(k)^{*}}$ with ${\rm flip}(a\otimes b)=b\otimes a$. Thus, the positions of $\alpha_{1}$ and $\alpha_{2}$ can be exchanged in the precedent proof, repeating mutatis mutandi, we reach that $T_{k}^{*}=T_{{\rm filp}(k)^{*}}\in S_{t}$, with $\frac{1}{t}=\frac{1}{2}+\frac{1}{p_{1}}$ and $\alpha_{1}>\frac{d}{p_{1}}$, this is a consequence of case (1) and by engaging the fact that $\|T_{k}\|_{S_{t}}=\|T_{k}^{*}\|_{S_{t}}$.
(3). If $\alpha_{1}=\alpha_{2}=0$, then the situation is trivial since $k\in L^{2}({\mathbb{T}}_{\theta}^{d}\otimes({\mathbb{T}}_{\theta}^{d})^{\rm op})$, it follows that
$T_{k}\in S_{2}\subset S_{r},\ r>2.$
Now the proof is complete. ∎
4. Acknowledgements
The authors were supported by the FWO Odysseus 1 grant G.0H94.18N: Analysis and Partial Differential Equations, the Methusalem programme of the Ghent University Special Research Fund (BOF) (Grant number 01M01021). Michael Ruzhansky is also supported by EPSRC grant EP/V005529/1 and FWO Senior Research Grant G022821N.